### v3

# What is glacier sliding?

Abstract
Glacier and ice-sheet motion is fundamental to glaciology. However, there is still no clear consensus for the optimal way to describe the sliding component of glacier and ice-sheet dynamics. Typically, sliding is parametrised using a traction coefficient linked to a given theory describing one or a limited set of sliding processes. However, this precludes the possibility of multiple overlapping and spatio-temporally varying sliding modes with inaccuracies resulting in model error propagation as the system evolves away from the conditions under which it was optimised for. Here, we describe glacier sliding as a scale- and setting-dependent ’inner flow’ that arises from multiple overlapping sub-processes, including viscous ice deformation and obstacle resistance (or ’form drag’). The corresponding ’outer flow’ then accounts for ice deformation that is only minimally influenced by bed properties. Following this framework, we suggest that a rough-smooth bed division may be more consequential than a hard-soft one, and that the importance of ’Iken’s bound’ can be significantly reduced if viscous ice deformation becomes a significant process in a given region within the inner flow, which may explain the persistent functionality of power-law sliding in large-scale process-agnostic sliding studies over rough topography. Further, reviewing observation-based studies and considering the diversity of sliding processes and settings, we suggest that a simple ’unified’ sliding relationship controlled by a single tunable coefficient may not be realistic. However, we show that given reasonable assumptions, sliding relationships should generally fall within a sum of regularised-Coulomb and power-law components, and suggest that a or compound, or power-law relationship with careful consideration of the power value may provide the most flexible way to account for the varied net effect of compound sliding sub-processes.
1 Background
The question then occurs, is the viscosity real or apparent? John Tyndall on glacier motion in 1857.
Glacier sliding is a phenomenologically distinct problem within earth sciences, separated from simpler plane-on-plane sliding problems between two elastic solids by multiple confounding factors. The relative softness of ice allows for both viscous deformation under a non-Newtonian power-law and discrete slip at the ice-bed interface, which may occur in a stick-slip or continuous manner. A ‘soft’ sediment phase provides a further medium to convey deformation in many instances (and is shaped in turn by glacial processes), while ‘hard’ underlying bedrock can constitute a highly topographically irregular base ranging from <1 cm asperities to dramatic fjords, necessitating ice deformation around rigid obstacles over length scales covering 5 orders of magnitude. A subglacial hydrological system comprised largely of melted ice may significantly alter the effective pressure of the ice upon its substrate, and influence the roughness and material characteristics of the base itself. Debris entrained within basal ice interacts with the subglacial environment and influences both basal ice properties and the frictional properties of the glacier sole. These considerations present a wide parameter space for basal boundary conditions, which may also exhibit considerable spatial and temporal variations. The net outcome is significant difficulty in the derivation of a ‘unified’ or ‘generalised’ sliding relationship, usually expressed in the form
$\tau_{b}=f(u_{b})\>\>\text{or}\>\>\tau_{b}=f(u_{b},N)$ (1)
where $\tau_{b}$ is basal traction, $u_{b}$ is ice velocity tangential to the ice-bed interface, $N=P_{i}-P_{w}$ is effective pressure, $P_{w}$ is subglacial water pressure, and $P_{i}$ is the ice overburden pressure (or with $u_{b}$ and $\tau_{b}$ vectorised for higher dimensions, e.g. Eq. 9).
Determining an appropriate sliding relationship is central to producing tractable, physically-based projections of the contribution of ice sheets and glaciers to sea-level rise and fresh-water fluxes over the coming centuries. Extant sliding relationships (Fig. 1, Table. 1) influence ice-sheet model output at catchment to entire ice sheet scales, with the distinction between bounded traction (plastic, regularised-Coulomb) and unbounded traction (all other relationships) being particularly consequential (Gillet-Chaulet and others,, 2012; Parizek and others,, 2013; Ritz and others,, 2015; Tsai and others,, 2015; Bons and others,, 2018; Kyrke-Smith and others,, 2018; Lippert and others,, 2024; Trevers and others,, 2024), resulting in up to a 100% increase in mass loss over 100 years dependent on initialisation (Brondex and others,, 2019; Åkesson and others,, 2021). Nonetheless, around five [^1] Some sliding relationships can be considered a simplified versions of one another. For example power-law and pseudo-plastic can be manipulated into linear or plastic relationships, the $N$ in Budd sliding is effectively subsumed into the $C$ of power-law sliding during inversion if it is assumed that subglacial water pressure does not vary in time, making this number subjective. (Section 3) sliding relationships implicitly or explicitly capturing different sliding processes are used in the Ice Sheet Model Intercomparison Project (ISMIP) models (Goelzer and others,, 2020; Seroussi and others,, 2020), which guide worldwide policy on sea level rise mitigation. This state of affairs is not immediately straightforward to resolve, as the inverse methods that tune for $C$, the ‘traction coefficient’, will produce reasonable basal stress states whichever equation in Table 1 is chosen as the global stress balance condition must be met (Joughin and others,, 2009; Arthern and Gudmundsson,, 2010). There is therefore no simple approach to discounting sliding relationships given solely a snapshot of glacier or ice sheet velocity and geometry. Errors associated with an inappropriate choice of basal sliding relationship subsequently increase as a function of model run time (Aschwanden and others,, 2021).
Figure 1: Existing sliding parameterisations and how their development relates to soft- and hard-bed theory. R-c+P-l represents Regularised-Coulomb plus power-law. Traction and velocity values are plausible, but for illustrative purposes only.
Sliding relationship Equation
Linear $\tau_{b}=Cu_{b}$  or  $\tau_{b}=NCu_{b}$
Plastic $\tau_{b}=C$  or  $\tau_{b}=CN$
Weertman(-type) $\tau_{b}=Cu_{b}^{\frac{2}{1+n}}$  or  $\tau_{b}=Cu_{b}^{1/m}$
Pseudo-plastic $\tau_{b}=\tau_{c}\frac{u^{1/m}}{u_{t}^{1/m}}$
Budd $\tau_{b}=C\left({N}^{q}u_{b}\right)^{1/m}$  or  $\tau_{b}=CNu_{b}^{1/m}$
Regularised-Coulomb $\tau_{b}=C\left(\frac{u_{b}}{u_{b}+u_{t}}\right)^{1/m}$
Table 1: Selected sliding relationships applied in numerical glacier and ice-sheet models, expressed in one-dimensional form. Denotes that the sliding relationship is used in Goelzer and others, (2020) or Seroussi and others, (2020) ISMIP experiments. In all equations $\tau_{b}$ is basal traction and $u_{b}$ is the basal ice velocity tangential to the bed. Where present, $n$ is the exponent used in Glen’s flow law (usually 3) and is only used where an explicit link to Glen’s flow law is made in the paper proposing the sliding relationship, $m$ is an exponent often related to Glen’s flow law, but not always explicitly, $N$ is the effective pressure, $C$ is the traction coefficient which may be adapted from its use in the original paper for simplified intercomparison (for example we replace $1/C$ with $C$ in the Budd relationship), $\tau_{c}$ is the yield stress in the pseudo-plastic relationship, $q$ is a fitting parameter in the Budd relationship, and $u_{t}$ is the threshold velocity in the pseudo-plastic and regularised-Coulomb relationships. The rate-weakening 2D cavities relationship in Fig. 1 is not included, as no numerical ice-sheet or glacier models include rate-weakening behaviour above a given velocity threshold. Depending on convention, the RHS may be negative in the original paper to indicate a traction force opposite to velocity direction. In dimensions higher than 1 the sliding relationship is vectorised (e.g. Eq. LABEL:eq:slipvel). Notes: Linear: Adapted from Nye, (1969) and Kamb, (1970). As applied in e.g. Morlighem and others, (2013). The version with $N$ is used in some Goelzer and others, (2020) experiments. Weertman: Equation featuring $n$ is adapted from Weertman, (1957) where $n$ is set as 3. $m$ is often used in models instead of $n$ where $m$ is usually between 3 and 4. Pseudo-plastic: Effectively the same as Weertman but included for frequent use in PISM applications e.g. Aschwanden and others, (2016), the pseudo-plastic relationship can be varied between plastic ($m=\infty$), linear ($m=1$), or power-law ($2\leq m\leq 4$) behaviour. Regularised-Coulomb: The simplest form excluding $N$, adapted from (Joughin and others,, 2019). Other formulations exist such as $\tau_{b}=Cu_{b}N\left(\frac{{u_{b}}^{-n+1}}{u_{b}+A_{s}C^{n}{N}^{n}}\right)^{1/n}$ (Helanow and others,, 2021) where $A_{s}$ is a parameter specific to this formulation. The differences between regularised-Coulomb relationships are not negligible (Fig. A1). Plastic: As used in e.g. Bougamont and others, (2014) and the PISM default (Winkelmann and others,, 2011). Budd: As developed in Budd and others, (1984) where $q=1$ and $m=3$ and used in Budd and Jacka, (1989) with $q=2$ and $m=1$. The alternative and simplified Budd formulation is as used in Choi and others, (2022). Tsai and others, (2021) also produces a relationship similar to Budd sliding.
In numerical ice-sheet models, these sliding relationships form a critical thread between the processes occurring at the sub-grid-scale and the numerical solution defined at the grid scale (where ‘grid’ refers to discretised model grid cells or mesh elements)  [^2] This was not always the case, with purely analytical sliding relationships predating the first numerical models by decades (Weertman,, 1957; Rasmussen and Campbell,, 1973; Jenssen,, 1977) which may explain why sub-grid-scale processes were not initially a concern.. In discretised continuum mechanics models, the distinction between sub-grid-scale and grid-scale is usually made using a ‘representative volume element’ that is large enough to be representative of the bulk behaviour of micro-scale processes within the constitutive material it represents. For ice deformation, this is straightforward – an ice volume of 1 m3, much below the resolution of even the most detailed numerical models, will contain sufficient variability in crystal orientation, size, and impurities as to have effectively the same response to an applied stress field as a nearby 1 m3 of the same material. Lab-based experiments using small volumes of ice (e.g. Glen,, 1952) can therefore be reasonably applied to much larger-scale models, even if separate issues such as the importance of tertiary creep remain (Adams and others,, 2021).
There is not, however, an obvious representative volume, or ‘representative surface’ element size for glacier sliding processes, with investigations into glacier sliding covering a range of scales extending over seven orders of magnitude from millimetres to 10s of kilometres (Fig. 2). For example, lab experiments for ice-sediment shearing behaviour may be conducted at a scale of tens of centimetres (e.g. Iverson and others,, 1998; Zoet and Iverson,, 2020), up-scaled to field settings covering an entire glacier where subglacially entrained clasts, variable subglacial hydrology, or bedrock obstacles further influence the relationship between traction and velocity (e.g. Iverson and others,, 1995 Hedfors and others,, 2003, Gimbert and others,, 2021), and ultimately applied to ice-sheet models where individual grid cells can exceed 10s of kilometres (or have a highly variable size across a given domain) and cover an uncertain and/or heterogeneous subglacial landscape (e.g. Kyrke-Smith and others,, 2018; Holschuh and others,, 2020; Paxman and others,, 2021). Nonetheless, routinely-used sliding parameterisations in ice sheet models almost exclusively use theories derived for one or a limited set of sliding processes at small scales (Goelzer and others,, 2020; Seroussi and others,, 2020), without representing the full range of scale-dependent processes.
Figure 2: Spatial scales for theories and parameterisations of basal sliding. A solid line indicates coverage as explicitly defined in the associated paper paper while a dashed line indicates probable situation dependent coverage and an arrow indicates extension beyond the scale bar or an uncertain coverage beyond the given limit. Reasoning behind the positioning of each spatial range is provided in Appendix A1.
Representation of sliding in glacier and ice-sheet models is further hindered by a tendency to classify the bed either as ‘soft’ (less rigid than ice and subject to sediment deformation and transport under glacier slip and subglacial hydrology even at short time scales) or ‘hard’ (more rigid than ice and only appreciably modified by sliding or hydrology at glacial-cycle time scales). Glacier and ice-sheet models generally cannot represent settings that may feature both soft- and hard-bed characteristics within a single grid cell or across a model domain (Fig. 3). This ‘one or the other’ classification is challenged by geological and geomorphological evidence of mixed soft- and hard-bed regions in both recently deglaciated landscapes (e.g. Hogan and others,, 2020; Garcia-Oteyza and others,, 2022) and in active subglacial settings (Jordan and others,, 2023) but has received very little attention in glaciological studies (Koellner and others,, 2019). A middle-ground between hard and soft beds can also exist if the substrate is only deformable under certain conditions, i.e. very thick ice cover with high subglacial water pressure. Sub-grid-scale topography is furthermore known to significantly influence ice flow (Kyrke-Smith and others,, 2018; Hoffman and others,, 2022; Law and others,, 2023), and may even dominate grid-scale basal traction (Fig. 3b), but is also lacking from present sliding parameterisations.
Figure 3: Typically sized grid cells and mesh element overlain on previously glaciated regions. Thwaites Glacier bathymetry is from Hogan and others, (2020). Frafjord topography is from Kartverket.no.
In this paper we provide a clear framework for describing glacier sliding as behaviour emerging from multiple scale- and setting-dependent processes (summarised in Fig. 4), rather than as a single dominant process. We reconsider the ‘inner-outer’ flow division (Section 2) first proposed by Fowler, (1977), but with a more universal approach to the inclusion of numerous sliding sub-processes over three dimensions and resistance from sub-grid-scale topography (Section 3). We briefly cover standard soft- and hard-bed theory (Sections 3.1 and 3.2); form drag, roughness, and temperate ice (Section 3.3); stick-slip vs. continuous sliding (Section 3.5); the overlap between soft- and hard-bed sliding (Section 3.6); the basal ice layer and ice rheology (Section 3.4); and how these sub-processes are modified by subglacial hydrology (Section 4). We focus on annual timescales as these are the most relevant for predictive and palaeo models, and largely avoid the transient influence of e.g. supraglacial lake drainages (Das and others,, 2008), rainfall events (Doyle and others,, 2015), or seasonal modulation (Bartholomew and others,, 2011; Sole and others,, 2013). In Section 5 we review process-agnostic studies fitting sliding relationships to observational data. Last, we consider if a simple (i.e. controlled by a single parameter) unified sliding relationship is a realistic possibility, consider the settings and scales under which Iken’s bound may not hold, and make suggestions for the application of sliding relationships to models (Section 6).
Throughout this paper we use sliding to refer to the net effect of processes that fall within the ‘inner flow’ (covered below), and slip to refer to discrete slip at the ice-bed interface where a clear ice-bed interface can be discerned. Unless otherwise specified, we refer to sliding situations where the ice-bed interface is at the pressure melting point. We hope that this paper is useful to those seeking an introduction to glacier sliding and coverage of advancements over the last decade, but we do not attempt an exhaustive review. Earlier reviews (Clarke,, 2005; Fowler,, 2010) and book chapters (Benn and Evans,, 2010; Cuffey and Paterson,, 2010) provide more comprehensive coverage. Our discussion is limited to continuum-scales in ice and we do not cover crystal-scale deformation or molecular-level interaction – Schulson and Duval, (2009) and Krim, (1996) provide more information on these topics.
Figure 4: Schematics of glacier sliding sub-processes and controls discussed in the text. Scale is intentionally omitted but scale generally increases left to right and top to bottom through sub-processes. Hydrological processes in k and l are considered as controls, since they do not contribute directly to ice transport; these are distinguished with blue outline boxes. a Form drag over sedimentary clasts. b Shear of sediment. c Ice with clasts sliding over a flat hard bed. d Stick-slip events. e Regelation. f Deformation within the basal boundary layer. g Cavitation. h Spatially variable sliding rates over rough topography. i Spatially variable ice deformation over rough topography. j Temperate layer processes. k Hydrology and channelisation. l Subglacial lakes.
2 Inner-outer flows
Fowler, (1977), Fowler and Larson, (1978), and Fowler, (1981) (hereafter Fowler1977) consider glacier sliding over a two-dimensional free-slip bumpy bed analytically. Recognising that the smooth basal boundary used in larger-scale analytical models is not a realistic representation of a rough glacier bed, Fowler1977 separate ice motion into an ‘inner’ flow and an ‘outer’ flow (reinterpreted in Fig. 5). The inner flow, which includes a component of basal slip, is theorised to closely follow the actual bed topography. The outer flow accounts for internal deformation at the typical depth and length scale of the entire glacier and is considered to be ‘sliding’ over a smoothed representation of the actual bed (i.e. the model bed). The outer and inner flows are then joined by a sliding relationship under the condition that the inner flow ‘feels’ the outer flow as a uniform shearing flow, and the outer flow ‘feels’ the inner flow as a tangential stress at the smoothed bed boundary.
Figure 5: Inner and outer flows. a Outer flow and inner flow with matching surface (dash-dot line) which may be similar, but is not directly equivalent, to the model bed surface. Depending on setting the inner flow may feature any of the processes in Fig. 4. b The inner flow is collapsed to occur at the model bed surface. The outer flow represents ice deformation that is not significantly influenced by bed processes.
In our application, the inner flow accounts for the relevant subset of processes outlined in Fig. 4 and Section 3 (as opposed to the simpler set of processes considered in Fowler1977) and the outer flow comprises englacial ice deformation with minimal influence from the processes in Fig. 4. In a modelling sense, the outer flow represents all motion not captured by the basal sliding component. A more complex example of multiple sub-processes operating in an inner flow is given schematically in Fig. A2.
An inner-outer flow distinction may facilitate a more holistic description of glacier sliding, but a consistent definition for the position of the inner-outer boundary does not immediately materialise – even if an explicit inner-outer boundary is in keeping with its implicit use in previous studies (e.g. Schoof and Clarke,, 2008; Zoet and Iverson,, 2020; Helanow and others,, 2021). The inner flow does not neatly fit the idea of a boundary layer as typically used in fluid dynamics to describe a region of turbulent flow separation (Batchelor,, 1967 and discussed further in Section 3.3) and shares only broad similarities with a bottom boundary layer as used in oceanography models (Campin and Goosse,, 1999). A numerical approach also means that the inner and outer flow can not be matched analytically as they are in Fowler1977. However, perturbations in strain rates tend to decrease in amplitude with increasing height above the bed, either exponentially (Balise and Raymond,, 1985) or as a power-law (Chandler and others,, 2006); corresponding length scales for the decay of inner layer strain rate perturbations could then form a basis for defining the inner-outer layer boundary.
The height at which inner-flow processes become negligible will also vary as a function of the spatial scale under consideration, heterogeneity in bed properties, and the roughness characteristics of the bed topography. For example, inner-flow processes will be important at a greater height above the bed if considering an area of rough topography around a planar soft-bed region, than if considering the planar soft-bed region in isolation. In most situations however, a reasonable norm is already defined for setting the position of the model bed surface. BedMachine (Morlighem and others,, 2017) and BedMap (Fretwell and others,, 2013) products are used for the Greenland and Antarctic ice sheets respectively while a more fragmented approach is taken for mountain glaciers (e.g. Welty and others,, 2020; Grab and others,, 2021). These bed topography models are significantly smoother than the roughness suggested by geostatistical analysis of radar flight lines (MacKie and others,, 2021; Law and others,, 2023) and deglaciated forelands immediately adjacent to the ice sheets, but function as convenient thresholds for separating inner and outer flows. And, even if there is full knowledge of the bed topography (for example modelling the palaeo-Scandinavian Ice Sheet), model resolution dictates some inherent reduction in fidelity of bed topography. Crucially, we note that the imperfect representation inherent to subglacial bed topography products means a degree of ice deformation will unavoidably already be incorporated within existing sliding parameterisations even if its inclusion has not been previously formalised. The error associated with a collapsed inner flow is therefore not introducing new errors, but rather providing a way to quantify and address existing ones.
Using this description of the inner flow, the net basal traction, $\tau_{b}$ or $\boldsymbol{\tau}_{b}=(\tau_{b,x},\tau_{b,y})$ for higher-dimensional settings, at the model bed surface can then be treated as the sum of multiple individual sliding processes and upslope bed resistance. Here, we separate these inner-flow processes into traction components for the following categories: resistance to discrete slip at the ice-bed interface represented as $\mathcal{S}$, resistance to distributed viscous deformation within the ice as as $\mathcal{D}$, and geometric resistance to flow from up-slope obstacles as $\mathcal{G}$, with each printed in bold font to represent vector components. In a one dimensional case we can relate $\mathcal{S}$ to the slip velocity at the ice-bed interface, $u_{s}$, $\mathcal{D}$ to the velocity gradient between the ice-bed interface and the inner-outer boundary, $u$, and $\mathcal{G}$ to the slope orientation, $\theta$, though we remain agonistic about the exact functions tying these terms together for now. Summing these components for a simple one-dimensional case gives
$\tau_{b}=\sum_{i}\mathcal{S}_{i}(u_{s})+\sum_{i}\mathcal{D}_{i}(u)+\mathcal{G}_{i}(\theta)$ (2)
for each sub-process $i$ within each category.
Considering next a region, $R$, that could be taken as a model cell (Fig. 6), and which may contain sliding processes that vary through space, where the plane defined by $x$ and $y$ follows the average of the plane defining the inner-outer boundary over the region which is also parallel to $\boldsymbol{\tau}_{b}$ and $\boldsymbol{u}_{b}$, the vector defining velocity at the inner-outer boundary. We can then define each sub-process as additionally being a function of space using an indicator function $I_{R_{i}}(x,y)$ dependent on a sub-region $R_{i}$ where
$I_{R_{i}}(x,y)=\begin{cases}1&\text{if }(x,y)\in R_{i}\,,\\ 0&\text{otherwise}\,.\end{cases}$ (3)
with each $R_{i}\subseteq R$. Focussing on traction, this gives a total, $\boldsymbol{\mathcal{S}}_{T}$, of slip processes as
$\boldsymbol{\mathcal{S}}_{T}(x,y,\boldsymbol{u}_{s})=\sum_{i\in\{\alpha,\beta,\gamma,...\}}I_{R_{i}}(x,y)\boldsymbol{\mathcal{S}}_{R_{i}}(\boldsymbol{u}_{s})\,$ (4)
where the subscript $R_{i}$ refers to a slip process occurring in $R_{\alpha}$, $R_{\beta}$ and so on (Fig. 6) and $\boldsymbol{u}_{s}=(u_{s,x},u_{s,y})$ refers to the slip velocity at the ice-bed interface within each $R_{i}$ which can be defined using local basis vectors. For practical applications to discretised grid cells $b_{io}$, the inner-outer boundary layer, should therefore define a plane. Appendix A3 shows an example of $\boldsymbol{\mathcal{S}}_{T}$ expanded into cases for three overlapping regions. This formulation furthermore follows most work on glacier sliding in assuming velocity is continuous in time (i.e. not stick-slip). Similarly for distributed processes
$\boldsymbol{\mathcal{D}}_{T}(x,y,\boldsymbol{u},b_{b},b_{io})=\sum_{i\in\{a,b,c,...\}}I_{R_{i}}(x,y)\boldsymbol{\mathcal{D}}_{R_{i}}(\boldsymbol{u},b_{b},b_{io})\,$ (5)
where $b_{b}$ is the position of the ice-bed interface, and $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ refers to the velocity field within the zone of deformation considered. Geometric resistance is then defined as
$\boldsymbol{\mathcal{G}}(F_{n},\boldsymbol{n})=-I_{F_{n}}(F_{n})\binom{F_{n}\cdot n_{x}}{F_{n}\cdot n_{y}}$ (6)
where $F_{n}$ is the normal force at a given point and $\boldsymbol{n}=(n_{x},n_{y},n_{z})$ is the corresponding unit normal vector defining the orientation of the ice-bed interface.$I_{F_{n}}(F_{n})$ is another indicator function defined as
$I_{F_{n}}(F_{n})=\begin{cases}1&\text{if }F_{n}\geq 0\,,\\ 0&\text{if }F_{n}<0\,.\end{cases}$ (7)
used in order to avoid integrating over regions of negative normal force. In practice and at longer (i.e. annual) time scales there should be no persistent negative normal force. If cavities (or subglacial lakes) are present there are two strategies to deal with them under this framework. First, we can assume the cavities are fully developed and take the water-ice interface where $F_{n}\simeq 0$ as the de facto ice-bed interface. The interface between ice and water can then simply be accommodated with a $\boldsymbol{\mathcal{S}}_{R_{i}}$ function (Eq. 4) through a frictionless relationship. Second, if $R$ covers a sufficiently large area, cavitation processes at a small scale can be subsumed within $\boldsymbol{\mathcal{S}}_{T}$ as a slip process (e.g. Helanow and others,, 2021). We are therefore not proposing a new way to simulate cavities, rather a new way to incorporate them into a larger framework.
$\boldsymbol{\mathcal{S}}_{T}$, $\boldsymbol{\mathcal{D}}_{T}$, and $\boldsymbol{\mathcal{G}}$ can then be integrated over the region $R$, with area $A_{R}$ to provide the area-averaged traction at the model bed interface
[…]
which could additionally be modified to include e.g. distributed deformation within a sediment layer through an additional term within the double integral.
Figure 6: Sub-regions related to sub-processes within a given region $R$. $R_{\alpha}$ and $R_{\beta}$ refer to slip sub-processes (Eq. 4), with the sub-region labelled $R_{\alpha}\cap R_{\beta}$ featuring both $R_{\alpha}$ and $R_{\beta}$. $R_{a}$, $R_{b}$, and $R_{c}$ refer to distributed deformation processes (Eq. 5) extend over height $h$. Dot-dash lines indicates the possible extension of region $R$ from a curved shape to a rectangular grid cell. We keep the vertical positions of $b_{b}$ and $b_{io}$ constant in this diagram, but this does not need to be the case.
It is further possible to remedy our agnosticism of the functions relating traction to velocity through the use of generic constitutive relationships for slip and deformation. Setting a given
$\boldsymbol{\mathcal{S}}_{R_{i}}=-I_{R_{i}}(x,y)P_{R_{i}}\left(\frac{|\boldsymbol{u}_{s}|}{|\boldsymbol{u}_{s}|+{u_{t}}_{R_{i}}}\right)^{\frac{1}{m_{R_{i}}}}\frac{\boldsymbol{u}_{s}}{|\boldsymbol{u}_{s}|}$ (9)
where $P_{R_{i}}$, and $m_{R_{i}}$, and ${u_{t}}_{R_{i}}$ are constants for a given region $R_{i}$ and which can represent the regularised-Coulomb relationship in Table 1, with the possibility to additionally incorporate $N$ if it is considered to be fully parametrised within $P$, or in a similar manner for a power-law relationship. A generic viscous distributed deformation sub-process can then be
[…]
where $b_{b}$ is the position of the ice-bed interface, $b_{io}$ is the height of the inner-outer flow boundary (Fig. 6), $\dot{\epsilon}_{x}^{2}=\dot{\epsilon}_{xi}\dot{\epsilon}_{xi}$ and $\dot{\epsilon}_{y}^{2}=\dot{\epsilon}_{yi}\dot{\epsilon}_{yi}$ are the effective strain rates in the $x$ and $y$ directions respectively, obtained from the strain rate tensor $\dot{\boldsymbol{\epsilon}_{ij}}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{j}}\right)$. Similarly to Eq. 9, $K_{R_{i}}$ and $n_{R_{i}}$ are constants for a given region $R_{i}$, but can be related to the parameters $A$ and $n$ respectively in the commonly used Nye-Glen isotropic flow law (Nye,, 1952; Glen,, 1955), $\dot{\boldsymbol{\epsilon}}=A\tau_{e}^{n-1}\boldsymbol{\tau}$, where $\tau_{e}^{2}=\frac{1}{2}\mathrm{tr}(\boldsymbol{\tau}^{2})$ is the effective stress and $\boldsymbol{\tau}$ is the deviatoric stress tensor.
It is more challenging to suggest a boilerplate equation for $\boldsymbol{\mathcal{G}}$. Lithostatic pressure will balance, and $P_{w}$ shouldn’t have a directional influence if we continue taking the ice-water interface as the de facto ice-bed interface, so dynamic pressure and surface slope will control the net effect of $F_{n}$ integrated over $R$, as studied by e.g. Weertman, (1957), Nye, (1970). At very low Reynolds numbers low-angled obstacles provide minimal or negligible resistance, but steep and abrupt obstacles may present important resistance contributions, which cannot be calculated analytically except in special cases (Fox and others,, 2006). Usually, the resistance, $F_{D}$, from a single more prominent obstacle is defined as
$F_{D}=C_{D}\frac{1}{2}\rho V^{2}A$ (11)
where $C_{D}$ is the drag coefficient specific to that obstacle and usually obtained experimentally, $V$ is the background velocity around the obstacle, and $A$ is the cross-sectional area of the obstacle opposing flow (Fox and others,, 2006, 9.7). Considering a sphere moving through a creep-dominated fluid with very low (<<1) Reynolds number, the scaling relationship is
$F_{D}\sim K_{R_{i}}V^{\frac{1}{n}}L^{2-\frac{1}{n}}$ (12)
(Appendix A2) where $L$ is the length scale of the obstacle. Therefore, while we stop short of suggesting an explicit relationship for $\boldsymbol{\mathcal{G}}$ here, given the uncertainty in its form for potentially complicated bed geometries, it is sufficient to say for now that $\boldsymbol{\mathcal{G}}$ will have negligible importance for very smooth beds, but be an increasing function of velocity, obstacle prominence, $\mathcal{O}_{p}$, and obstacle frequency, $\mathcal{O}_{p}$, and which may follow a power law similar to $\boldsymbol{\mathcal{D}}_{R_{i}}$. We represented this functionally as
$\boldsymbol{\mathcal{G}}=f(\boldsymbol{u}_{b},\mathcal{O}_{p},\mathcal{O}_{f})\,.$ (13)
We leave it to future studies to better quantify $\boldsymbol{\mathcal{G}}$ in the context of realistically rough landscapes, but provide Eq. 13 as scaffolding in the meantime.
Overall, this framework extends from Fowler1977 to include arbitrarily many sub-processes in three dimensions and facilitates quantification of the role of each sub-process in the inner flow, though a full representation will be further complicated in reality; for example if we consider multiple rheological layers between $z=b_{b}$ and $z=b_{io}$. Form drag is covered in greater detail later (Section 3.3), but we henceforth intend the sum of $\boldsymbol{\mathcal{D}}$ and $\boldsymbol{\mathcal{G}}$ when we refer to form drag processes. We consider how this framework could relate to a unified sliding relationship in Section 6.3.
3 Sliding processes
3.1 Sliding over ‘soft’ (sediment) beds
Early work on soft-bed sliding suggested that subglacial till behaves as a mildly non-linear Bingham viscous fluid (such that resistance to deformation increases with deformation rate), with distributed deformation extending around 50 cm below the ice-bed interface (Boulton and Hindmarsh,, 1987). This idea received some support from contemporaneous studies of subglacial conditions (Blake,, 1992; Humphrey and others,, 1993) but most subsequent studies – including laboratory, direct measurement, and analogous laboratory experiments from different material science disciplines – find strain-localisation close to the ice-bed interface, or manifest in multiple discrete shear zones at multiple depths within the subglacial till, and a Coulomb-plastic rheology (e.g. Biegel and others,, 1989; Kamb,, 1991; Hooke and others,, 1997; Iverson and others,, 1997; Iverson and others,, 1998; Iverson and Iverson,, 2001; Damsgaard and others,, 2013; Cuffey and Paterson,, 2010), and we therefore do not discuss viscous sediment deformation further.
In laboratory experiments of glacial-till deformation beneath a rigid shear ring, the till reaches its yield strength and then fails at a uniform rate under steady conditions, with a near-constant or a modest decrease in basal traction over a large (400 m a-1) shear-ring velocity increase (Iverson and others,, 1998; Fischer and others,, 2001; Iverson,, 2010). Field observations using strain-gauges emplaced in actively deforming subglacial sediment repeatedly support Coulomb-plastic deformation (Iverson and others,, 1995; Alley and Whillans,, 1984; Fischer and others,, 2001), with an inverse relationship between pore-water pressure and sediment strength (Fischer and others,, 2001). Numerical models indicate that the pore-water pressure-strength relationship is further modulated by till permeability properties (Damsgaard and others,, 2017, 2020), with till properties influenced in turn by crushing and compression from overlying ice (Iverson,, 1999). Field observations additionally emphasise the importance of ploughing of clasts lodged at the ice-till interface, particularly where sediment pore pressure is high, as the pressure exerted by the clast on the downflow sediment can locally increase the water pressure and weaken the sediment (Iverson and others,, 1994; Rousselot and Fischer,, 2005). In the case where sediment is not weakened by an increase in pore pressure, a clast may be pushed downwards and no longer be ploughed through the matrix (Clark and Hansel,, 1989).
More recently, Zoet and Iverson, (2020) used a ring of temperate ice $\sim$15 cm thick between the rigid shear ring and underlying sediment. In this setting, rate-strengthening behaviour occurs at lower velocities (below around 50 m a-1 in sediment with clasts, Fig. 4a) as ice viscously deforms and regelates around static clasts before the clasts begin to plough through the finer-grained matrix at a stress limited by the till’s Coulomb strength (Fig. 4b). The net result of the Zoet and Iverson, (2020) experiment is a regularised-Coulomb relationship (Fig. 1, Table 1) where the sliding interface is defined up 15 cm above the ice-till interface.
Bedforms, including drumlins, mega-scale glacial lineations, and ribbed moraine (Stokes,, 2018) are ubiquitous in areas of continuous sediment and likely represent a continuum of features formed by glacier dynamics and subglacial hydrology (Ely and others,, 2023). These bedforms are generally streamlined and have a low aspect ratio (height/length), but their contribution to form drag has not been quantified to our knowledge.
3.2 Sliding over ‘hard’ (bedrock) beds
Hard-bed sliding studies can be divided into those considering slip across planar surfaces of limited extent and those concerned with sliding over rough or undulating beds at longer wavelengths. In studies considering planar surfaces water pressure is typically kept spatially constant, while in studies considering a rough or undulating bed subglacial water may be theorised as either a microscopically thin and continuous layer, or as a pressurised system of subglacial cavities occupying local bedrock depressions. Undulating beds with a frictionless ice-bed interface were the first to attract attention in sliding theory, but we begin with macroscopically flat (i.e. flat to the naked eye) surfaces due to their simplicity. Roughness at length scales beyond 25 m is covered in Section 3.3.
The slip of ice over a planar surface (Fig. 4c) is closer to the assumptions of standard Coulomb behaviour (Coulomb,, 1785; Desplanques,, 2014), though plane-normal velocity components, englacial clasts, and possible clast-scale cavitation present complications (Hoffman and others,, 2022). Rather than a sliding relationship of the form of Eq. 1, basal traction in these studies is often reported as a static $\mu_{s}$ or transient $\mu_{t}$ coefficient of friction, $\mu_{s}\>\textup{or}\>\mu_{t}=\frac{\tau_{b}}{N}$ where $\mu_{s}$ describes the force required to move the plane from stationary and $\mu_{t}$ describes the force required to keep the plane in motion. Shear-ring studies using temperate ice and englacial sediment content of 0-20% over macroscopically flat granite find a $\mu_{s}$ of 0.02-0.05 (Barnes and others,, 1971; Zoet and others,, 2013; McCarthy and others,, 2017; Thompson,, 2020), lower than teflon-on-teflon with no lubrication (Fetfatsidis and others,, 2013). In the above studies, sliding may be stick-slip or steady-state (outlined further in Section 3.5) depending on water pressure and particular material properties. Transient coefficients of friction between 0.05-0.08 were obtained from subglacial access to the ice-rock interface of the Engabreen outlet glacier in Norway (Iverson and others,, 2003; Cohen and others,, 2005). Unexpectedly high shear traction values of up to 0.5 MPa accompany the low $\mu_{t}$ values at Engabreen and these are still not fully explained, suggesting a possible gap in understanding – see Thompson and others, (2020) for a more in depth discussion (and note that, in calculations of transient friction, in contrast to static friction, the line defined by $\mu_{t}$ will not necessarily pass through the graph origin of $N$ and $\tau_{b}$). The transient coefficient of friction is much higher for cold ice (0.5 at -20oC) making sub-temperate slip very low, but potentially non-negligible (Barnes and others,, 1971; McCarthy and others,, 2017; Atkins,, 2013; Mantelli and others,, 2019; Section 3.7).
Where an ice block rests unconstrained on a plane, and water pressure beneath the ice is close to atmospheric pressure, $\mu_{s}$ can reach 0.6 for a rough (up to 0.25 cm roughness) pebbly surface or 0.2 for smooth concrete (Budd and others,, 1979). These are not typical stress or hydrological conditions for glacier beds (e.g. Hubbard and Nienow,, 1997; Woodard and others,, 2021) but were used by Budd and others, (1979) to derive an empirical sliding relationship (Table 1), with later modifications to exponents to produce a better fit with field data at an ice-sheet scale (Budd and others,, 1984) that are arguably large enough to mark a departure from the empirical underpinnings. To summarise planar hard-bed sliding, it is reasonable to treat the friction of ice on rock in the presence of pressurised water as low, but it is not negligible, and may in certain configurations be an important local control on basal traction. There is significant scope to improve understanding in the processes responsible for ice-rock friction across a wide parameter space (Thompson and others,, 2020).
Sliding over rough hard beds has also received much attention. This began with the classical work of Weertman, (1957), who considered ice motion over isolated cuboids (akin to the geometry in Fig. 4e) with no ice-bed separation through two processes: regelation and enhanced creep (Fig. 4e). In regelation, ice moves without deformation. Temperate ice approaching an obstacle is subject to greater pressure, lowering its melting point. Consequently, some of the ice melts at the upstream face, and the meltwater is driven to the downstream face by the pressure gradient. Here, where the pressure is lower and the melting point is higher, the meltwater refreezes. The higher temperature at the downflow side of the obstacle results in heat flow back through the obstacle, completing the cycle. Regelation is an ongoing process in many subglacial settings (Kamb and Lachapelle,, 1964; Kamb,, 1970; Hallet and others,, 1978; Hubbard and Sharp,, 1993; Iverson,, 2000; Cook and others,, 2011; Rempel and Meyer,, 2019), including in soft bed settings outside the scope of the original Weertman, (1957) paper. However, the requirement for return heat flow causes regelation rate to vary inversely with obstacle wavelength, limiting the importance of regelation as a sliding process to obstacles smaller than a few cm (Weertman,, 1957). Regelation is therefore generally not viewed as a significant contributor to overall sliding rates (MacAyeal,, 2019) and has received comparatively little attention since the mid 1990s, though further contemporary study would not be unwelcome.
Enhanced creep is prompted by stress concentrations at bedrock obstacles, leading to locally increased strain rates. Functionally, this means enhanced creep can be considered as a synonym for form drag (Section 3.3), and would fall under $\boldsymbol{\mathcal{D}}_{T}$ (Eq. 5) in our framework (Section 2). Conversely to regelation, the contribution of enhanced creep to glacier sliding in Weertman’s 1957 study increases with obstacle size. Focusing on the obstacle sizes that produce the greatest rate of sliding when regelation and enhanced creep act in unison (millimetre- to metre-scale) Weertman, (1957) reaches a value of $m$ in a power-law of $(n+1)/2=2$ with $C$ representing a combination of bed-geometry characteristics. Technically, values of $m$ deviating from 2 when $n=3$ are therefore not physically based on the original Weertman theory making ‘power-law’, which we use henceforth, or ‘Weertman-type’ sliding a more suitable descriptor. Sliding over a sinusoidal two-dimensional bed with no ice-bed separation and a linear ice rheology produces a linear sliding relationship ($m=1$ in power-law sliding) (Nye,, 1969), as does Weertman sliding with $n=1$.
power-law sliding, usually with $m=3$, is the most commonly used sliding relationship in the latest round of ISMIP experiments (with linear sliding also frequently used) (Goelzer and others,, 2020; Seroussi and others,, 2020), yet the theoretical underpinnings supporting these relationships in the original papers do not hold up to scrutiny – including by Weertman’s own reappraisal (Weertman,, 1979). Weertman’s original theory features a non-linear rheology following Glen, (1955), but it is apparent that no glacier beds are actually characterised by a controlling obstacle size with the morphology used by Weertman, (1957), and even changes to a different unrealistic bed geometry produces a materially different relationship (Lliboutry,, 1979). Weertman sliding is further challenged by ubiquitous evidence and theoretical support for subglacial cavitation, or regions of bounded ponded water at high pressure between the glacier sole and underlying hard or soft bed, which form when the pressure on the downstream face of an obstacle falls below a critical level (Fig. 4g and e.g. Walder and Hallet,, 1979; Kamb,, 1987; Hooke,, 1989; Helanow and others,, 2021). Weertman sliding also does not provide an explanation for increased sliding rates when subglacial water pressure increases. The persistent utility of power-law sliding in heuristic studies, though generally with $m\geq 3$, despite these major problems is covered in Sections 5 and 6.3.
Subglacial cavities (Fig. 4g) act to reduce contact between the glacier sole and the ice bed and may grow as a result of increasing basal velocity or subglacial water pressure. In the two-dimensional formulations of Lliboutry, (1968), Iken, (1981), Schoof, (2005), and Gagliardini and others, (2007), amongst others, where the bed is comprised of frictionless sinusoids, this produces an upper bound for basal traction, followed by rate-weakening behaviour (Fig. 1). Put mathematically, ‘Iken’s bound’ as coined by Schoof, (2005) is given as
$\tau_{b}\leq N\textup{tan}(\theta)$ (14)
where $\theta$ is the maximum up-slope angle between the bed and the mean flow direction of the ice (Iken,, 1981; Schoof,, 2005). Roldan-Blasco and others, (2022) suggest that in the simplified two-dimensional framework of compound sinusoids non-negligible friction at the ice-bed interface will not significantly alter the form of the sliding relationship or the existence of Iken’s bound. However, the geometry used in these studies can also be considered unrealistic, with the absence of a third dimension preventing lateral escape of subglacial water which would (1) act to reduce the cavity size (Gimbert and others,, 2021) and (2) may place a limit on the maximum obtainable subglacial water pressure averaged over a large region at long (i.e. annual) time scales after the system has had time to equilibrate (e.g. Dow and others,, 2015; De Fleurian and others,, 2018; Hart and others,, 2022). Helanow and others, (2019) and Helanow and others, (2021) address the three-dimensional problem for smaller scales, and show that realistic three-dimensional bed topographies up to 25 m still result in bounded basal traction, but that significant rate-weakening behaviour is not expected (the regularised-Coulomb relationship in Fig. 1). The validity of the assumptions underlying Eq. 14 when considering inner flows over larger areas and rough topography is covered in Section 6.2.
3.3 Roughness, form drag, and temperate ice
Glaciated and deglaciated landscapes are characterised by roughness not only at scales of centimetres to metres (as featured in the previous section), but also by roughness at scales of tens to thousands of metres. This includes cnoc-and-lochan landscapes, drumlins, incised fjords, and other erosional or depositional features that bridge a very subjective boundary between ‘roughness’ and ‘topography’ (Figs. 4h-j). Radar flight lines reveal extensive and variably rough beds beneath the Greenland and Antarctic ice sheets (e.g. Bingham and Siegert,, 2009; Rippin,, 2013; Munevar Garcia and others,, 2023), often quantified through a Fast Fourier Transform or variogram (e.g. MacKie and others,, 2021) which pares back the geomorphological uniqueness of a landscape but enables regional intercomparison. At present, only roughness up to a length scale of 25 m is explicitly incorporated in a process-based sliding relationship (Helanow and others,, 2021), but it is clear from borehole observations (Ryser and others,, 2014; Doyle and others,, 2018; Maier and others,, 2019; Law and others,, 2021) and simulations of ice motion over rough terrain at multi-kilometre scale (Hoffman and others,, 2022; Law and others,, 2023; Liu and others,, 2023) that roughness at larger scales significantly influences patterns of ice motion. Deformation rates are highly variable in proximity to the bed (Fig. 4i) with basal slip ranging between 5%-95% of total surface velocity (Fig. 4h) and much higher slip rates over topographic highs (Law and others,, 2023). Existing soft- and hard-bed sliding theories and parameterisations do not incorporate these processes – yet the limited studies that have investigated its influence on bed properties suggest a notable control (Wilkens and others,, 2015; Gillet-Chaulet and others,, 2016; Falcini and others,, 2018).
In polythermal glacier and ice-sheet settings, a layer of lower-viscosity basal temperate ice (Fig. 4j) may furthermore modulate the complex motion patterns caused by topographic roughness (Krabbendam,, 2016; Law and others,, 2023), with greater thicknesses of temperate ice found in topographic troughs than peaks, and the formation of a shear band at the top of the temperate zone (Law and others,, 2023; Liu and others,, 2023). The rheology of pure temperate ice has recently been constrained as likely linear viscous (Lliboutry,, 1971; Adams and others,, 2021; Schohn and others,, 2025), though the rheology of temperate ice with a high impurity content remains unconstrained through laboratory experiments to our knowledge.
The influence of roughness on sliding can be theorised through form drag, even if it is at fist glance complicated to delineate the transition between ice deformation contributing to presently-derived hard bed sliding relationships (Section 3.2), and ice deformation contributing to form drag. Kyrke-Smith and others, (2018) consider the influence of increasing the spatial resolution of a numerical model from 5 km to 0.5 km (and therefore also increasing the fidelity of the underlying topography) on the basal traction obtained from an inversion procedure. Intuitively, Kyrke-Smith and others, (2018) find that as spatial resolution is decreased, basal traction increases. Form drag is then described as the area-averaged basal traction with high topographic fidelity subtracted from the area-averaged basal traction with low topographic fidelity, i.e. the basal traction arising from topographic obstacles not explicitly represented by the basal boundary position. This clearly aligns with the idea that form drag can be theorised as part of the inner-flow which collapses to the modelled bed surface (Section 2). At a much smaller scale, Minchew and Joughin, (2020b) comment on Zoet and Iverson, (2020) and invoke form drag as the viscous resistance that arises as ice deforms over entrained clast ‘micro-topography’ in laboratory soft-bed sliding experiments before the sediment reaches its yield strength described previously in Section 3.1.
Form drag may therefore be a very useful concept in descriptions of glacier sliding, but its use in glaciology differs fundamentally from form drag as originally described for many aerodynamic applications. The differences are sufficiently large that it may be more appropriate to use a different phrase altogether but we maintain the the phrase ‘form drag’ for continuity with previous studies. In aerodynamic applications, form drag refers to the force acting on a solid body moving through a viscous and locally-turbulent fluid that is opposite and parallel to its velocity and which is not accounted for by induced drag, where the induced drag is a force resulting from turbulent vorticies connected to the lift on the body (Batchelor,, 1967, sections 5.11, 7.8). Form drag has also been used in atmospheric and ocean dynamics, where it typically refers to the entire drag force resulting from turbulence around an obstacle or obstacles fixed to the surface, rather than a division of it (e.g. Arya,, 1973; Renfrew and others,, 2019; Jagannathan and others,, 2023). Difficulties are still encountered when applying the concept of form drag to settings where there is a fixed continuous surface however, even if turbulence is present (MacCready and others,, 2003).
To circumvent the problem of defining form drag only at a specific scale – i.e. for ‘topographic’ roughness (Kyrke-Smith and others,, 2018) or for ‘sedimentary-clast’ roughness (Minchew and Joughin,, 2020a) – we define form drag here as $\boldsymbol{\mathcal{D}}_{T}+\boldsymbol{\mathcal{G}}$: the sum of all distributed viscous deformation and geometric contributions to the resistive stress operating over a given region within the inner layer. This means deformation processes across a very wide spatial scale (Fig. 2) including temperate ice (this section), basal ice layer processes (Section 3.4), and ’enhanced creep’ described in the original Weertman, (1957) paper are all taken as form drag components under our definition. We make this grouping lacking a consistent rule to meaningfully make distinctions between these processes.
3.4 The basal ice layer and ice rheology
A basal ice band of a distinctly different nature to clean meteoric ice, referred to here as the basal ice layer (Figs. 4f, 7), but which may also be called the frozen fringe (Meyer and others,, 2018), is frequently a feature of the lowermost section of glaciers and ice sheets. The basal ice layer is characterised by entrained debris and the diagenetic modification (or metamorphism) of meteoric ice by hydrologic processes, melting, refreezing, and intense strain (see Hubbard and Sharp,, 1989; Knight,, 1997; Souchez and others,, 2000; and Hubbard and others,, 2009 for comprehensive reviews). A basal ice layer is observed in deep ice-divide Greenlandic and Antarctic ice sheet ice cores (e.g. Gow and others,, 1979; Gow and Meese,, 1996; Souchez and others,, 1998; Tison and others,, 1998; Souchez and others,, 2002), Greenlandic and Antarctic ice sheet margins (e.g. Swinzow,, 1962; Tison and others,, 1993), and across many alpine and ice-cap settings (e.g. Hubbard and Sharp,, 1995; Lawson and Kulla,, 1978). Proposed formation mechanisms for the basal ice layer are basal freeze on entraining sediment through regelation (at a scale of up to around one meter) (e.g. Weertman,, 1961; Hubbard and Sharp,, 1993), a pressure-driven heat pump of temperate ice (up to a scale of several meters) (Robin,, 1976), or movement of subglacial meltwater along hydro-potential pathways to a region of the glacier bed below the pressure melting point (e.g. Bell and others,, 2011). Basal melting associated with frictional heat generated through slip is expected to decrease the thickness of the basal ice layer and impede its growth (Hubbard and Sharp,, 1989; Knight,, 1997). However, large basal ice layer thicknesses (up to 30 m) are recorded at the western terminus of the GrIS (Fig. 6, Knight and others,, 2002) and it is not conclusive if these sequences formed from freeze on in the immediate vicinity of the margin, or survived passage through up to 100 km of thawed bed conditions and slip-driven basal melt (MacGregor and others,, 2016). Recent modelling and radar mapping of Antarctica suggests thick (>100 m) sequences of sediment-laden basal ice may persist for tens of kilometres at least beyond a transition from frozen to thawed bed settings (Franke and others,, 2024) but further work is required to fully explore the parameter space of these processes. Further, spatially and temporally varying basal ice characteristics ranging from debris-rich ice to frozen sediment, will blur the distinction between soft- and hard-bedded regions.
Figure 7: The basal ice layer at Russel Glacier, Greenland. From Knight and others, (2002) showing characteristic stratigraphy.
Shear-ring laboratory experiments for the deformation of a 1-2 cm thick ice-sediment melange between thawed sediments and sediment (Hansen and others,, 2024) suggest similar regularised-Coulomb behaviour to shear-ring experiments of clean ice over sediment (Zoet and Iverson,, 2020). Meyer and others, (2018) also construct an analytical model arguing that ice infiltration into previously thawed sediments limits bed strength (in effect, plastic sliding) in pervasively soft-bedded regions. Beyond these studies however, we are not aware of experiments considering basal ice sequences exceeding a few cm in thickness or the large diversity of facies found in the basal ice layer (Hubbard and others,, 2009).
Four further rheological factors present significant complications for ice deformation in the vicinity of the bed and are hence captured at least partially within the inner flow. First, pre-Holocene ice with a higher dust concentration and typically smaller grain size deposited during the last glacial period deforms at a rate around 2.5 times that of Holocene ice under the same stress and temperature conditions (Paterson,, 1991). Such ice is widely present in the strata of the Greenland and Antarctic Ice Sheets (Macgregor and others,, 2015; Winter and others,, 2019; Ashmore and others,, 2020) but uncommon in glaciers and ice caps where it is limited to very slow moving areas (Thompson and others,, 1997), though valley glaciers generally feature much higher rates of englacial debris impurities in any case (e.g. Goodsell and others,, 2005). Second, ice at depth has a highly anisotropic rheology due to the development of a strong crystallographic preferred orientation under consistent uni-directional deformation (Lile,, 1978; Baker,, 1981; Duval,, 1981; Wilson and Sim,, 2002). Such mechanical anisotropy presents major challenges in implementation, bench-marking, and interpretation within numerical modelling (Martín and others,, 2004; Gillet-Chaulet and others,, 2005) but is nonetheless likely of significant importance within viscous deformation of ice in the inner flow where strain rates are both high and spatially complex (Law and others,, 2023). To our knowledge, this aspect of glacier sliding is not covered by any existing glaciological literature. Third, folds close to the bed may arise from mechanical differences in ice layers (Whillans and Jezek,, 1987; Bons and others,, 2016; Zhang and others,, 2023) with the possibility to significantly influence the bulk rheological behaviour of the inner flow. Fourth, while $n=3$ is the default in glacier and ice-sheet models (in fact already rounded down from the 3.2 reported in Glen,, 1955), there are numerous field-observation based studies suggesting that for cold (i.e. not temperate) ice $n=4$, representing a dislocation creep regime, is more appropriate in higher stress settings (away from ice-sheet ice divides, for example) (Bons and others,, 2018; Millstein and others,, 2022; Ranganathan and Minchew,, 2024) and perhaps also at areas dominated by pure shear (Gillet-Chaulet and others,, 2011). Considering a smooth BedMachine lower boundary in northern Greenland, $n=4$ dramatically decreases the area over which basal sliding is expected to contribute significantly to total surface displacement (Bons and others,, 2018). The influence of $n=4$ on form drag and sliding parameterisations, and its interaction with rough topography is to our knowledge entirely unexplored.
Where present, these rheological complications will already be implicitly and unavoidably accounted for via inversion procedures under parameterisations where they are not explicitly accounted for. The potential heterogeneity of their interactions creates significant complexity that future work should seek to disentangle.
3.5 Stick-slip and continuous sliding
In most cases, glacier and ice-sheet models neglect acceleration in the Navier-Stokes equations and treat ice as inelastic. These generally reasonable assumptions at typical modelling time-scales require that basal traction and basal velocity are connected through a continuous relationship in the form of Eq. 1 and that rapid (e.g. order of seconds) changes in velocity are not accounted for. However, these requirements are sometimes inconsistent with field observations of sliding, where near-instantaneous stick-slip behaviour (Fig. 4d) is frequently recorded by seismometers and geophones as basal icequakes (e.g. Neave and Savage,, 1970; Walter and others,, 2013) and also by direct observation at the glacier bed (Theakstone,, 1967; Hubbard,, 2002). Icequakes are also generated englacially by processes such as extensional faulting near the surface, but these phenomena are not covered here.
Basal icequakes occur due to a rapid release of elastically stored energy, resulting in ice motion that is transiently far higher than the long-term average (Winberry and others,, 2009; Podolskiy and Walter,, 2016). Moment magnitudes have been recorded from negative in alpine settings (Helmstetter and others,, 2015) to magnitude seven at Whillans Ice Stream in West Antarctica (Wiens and others,, 2008). For a basal icequake to occur, the local frictional shear strength of the bed must be lower when the sliding interface is in motion than when it is static, i.e. $\mu_{t}$ must be lower than $\mu_{s}$ (Bahr and Rundle,, 1996; Rice and others,, 2001) describing slip-weakening behaviour. Stick-slip is usually associated with diurnal or tidal variations in subglacial water pressure allowing a clear build-up period for elastic energy (Bahr and Rundle,, 1996; Bindschadler and others,, 2003; Walter and others,, 2008; Stevens and others,, 2024) but basal icequakes are also recorded outside of such cycles (Hubbard,, 2002). Stick-slip behaviour is usually described through a rate-and-state framework (Rice and others,, 2001) – as often applied in geological-fault settings (Gomberg and others,, 2000; Appendix A4) – where a state variable, $\psi$, is introduced to account for the strength of the fault (Rice and others,, 2001; van den Ende and others,, 2018), usually based on the slip displacement (Ruina,, 1983) or its time-dependent evolution (Dieterich,, 1979) giving
$\tau_{b}=f(u_{b},N,\psi)\>.$ (15)
In contrast to Eq. 1, the expectation is that $u_{b}$ in Eq. 15 can vary rapidly in both time and space. A rate-and-state friction relationship can still account for the aseismic rate-strengthening steady slip of a fault (van den Ende and others,, 2018) but is not well-suited for models entirely excluding rapid velocity changes and elastic behaviour by design.
Quantifying the contribution of stick-slip basal sliding to glacier motion is challenging due to large uncertainties in fault dimensions, slip displacement, and stress calculated from seismic data (Abercrombie,, 2015). Some estimates suggest stick-slip displacement accounts for all basal sliding over an annual period (Helmstetter and others,, 2015), while in other cases, coseismic and aseismic regions coexist (Barcheck and others,, 2020; Kufner and others,, 2021) – analogous to coexisting seismogenic and creeping faults in actively deforming geological faults (e.g. Azzaro and others,, 2020).
Despite the growing interest in stick-slip motion in glaciers (e.g. Podolskiy and Walter,, 2016; Zoet and others,, 2020) there are limited connections between rate-and-state theory (of form Eq. 15) and continuous basal traction theory relationships (of form Eq. 1). We are only aware of the statistical mechanical approach of Bahr and Rundle, (1995), where the glacier sole is represented as numerous elastically interconnected blocks with stochastically varying roughness parameters and the similar but simplified 1D approach of (Köpfli and others,, 2022). Bahr and Rundle’s approach, intended for thin, steep, valley glaciers, finds that the average total force on an individual block (including elastic forces from adjacent blocks) is proportional to the area-average basal traction, $\tau_{b}$. If this holds for greater ice thicknesses and larger scales then stick-slip behaviour alone will not invalidate continuous basal traction relationships over a sufficiently large grid cell (or representative surface area), but much further work is required to integrate this theory with recent developments in understanding the stick-slip behaviour of glaciers, and how this can relate to the spatial scales of numerical model grid cell resolution.
3.6 Overlap between soft- and hard-bed sliding
Most glacier and ice sheet modelling studies use basal traction relationships that implicitly assume the bed is either soft or hard across the entire domain with the ostensibly hard-bed Weertman-type sliding relationship (or equivalently, power-law/pseudo-plastic sliding with $2\geq m\geq 4$) being the modal choice in ISMIP6 experiments. In some situations a hard-soft division is appropriate, for example most studies are in agreement that extensive, relatively planar sediment lies beneath the ice streams draining into the Ronne and Ross ice shelves in Antarctica (Tulaczyk and others,, 2000; Vaughan and others,, 2003; Peters and others,, 2007), while geographically limited areas of Canadian continental shield may qualify as ‘true’, relatively planar, sediment-free hard beds (Slaymaker and Kovanen,, 2017). Many individual valley glaciers can also be reasonably categorised as predominantly soft- (e.g. Murray and Porter,, 2001) or hard-bedded (e.g. Hubbard,, 2002) (Section LABEL:s:glaciericesheet). However, taking Isunnguata Sermia in west Greenland as one well-studied example, borehole data mostly from topographic highs are used to suggest hard-bed conditions are dominant (Harper and others,, 2017; Maier and others,, 2019) while nearby seismic surveys are used to suggest the opposite (Booth and others,, 2012; Dow and others,, 2013; Kulessa and others,, 2017). Meanwhile, ice-marginal studies in west Greenland (Klint and others,, 2013) find a complex mix of bedrock, sediment-filled depressions, and extensive sandurs filling valley bottoms in front of the major land-terminating outlets (Grocott and McCaffrey,, 2017). In the case of the seismic survey of Kulessa and others, (2017) the recorded seismograms also allow for hard-bed conditions at topographic highs which would be in agreement with Harper and others, (2017). Similarly, recent work at Thwaites Glacier in Antarctica further suggests the coexistence of soft- and hard-bed regions over distances less than 10 km (Jordan and others,, 2023), supported by distinct along-flow variations in roughness (Holschuh and others,, 2020). While soft- or hard-bed categories can be a reasonable distinction for individual glaciers, all available evidence indicates that a more appropriate default for present and palaeo ice-sheets at large is a mixed bed condition vacillating between the two end-members.
Some recent attention has been focused on the interaction between soft- and hard-beds, with Koellner and others, (2019) considering 2D flow over sinusoidal beds with wavelengths from 8-60 km and setting the value of $m$ in the power-law sliding relationship (Table 1) to 8 to approximate regularised-Coulomb sliding in sinusoid troughs, while keeping it at 3 over the sinusoid peaks. This demonstrates that a hybrid $m=3\leftrightarrow 8$ Weertman rheology results in a modelled glacier response between $m=3$ and $m=8$ end-members. Given sliding rates at the modelled ice-bed interface can vary significantly over much shorter distances (Law and others,, 2023) than considered in Koellner and others, (2019), further work is needed to effectively isolate the influence of realistic variations in soft- and hard-bed sliding, and to assess whether hard-bed highs and soft-bed depressions are a reasonable model for the geologic composition of the Greenland and Antarctic ice sheets. Joughin and others, (2010) also apply a plastic sliding relationship to inferred soft-bedded regions and power-law sliding with $m=3$ to inferred hard-bedded regions Joughin and others, (2009) at Pine Island Glacier, Antarctica, finding good agreement in this mixed model over an 8-year period (Section 5).
From empirical studies at small (<25 m) scales, both the derived sliding relationship for hard-bed cavitation (Schoof,, 2005; Gagliardini and others,, 2007; Helanow and others,, 2021) and slip over sediments (Zoet and Iverson,, 2020) can yield a bounded traction relationship, leading to suggestions that regularised-Coulomb then represents a ‘universal glacier slip law’ for both hard and soft beds (Minchew and Joughin,, 2020b). We cover large-scale remote-sensing/modelling studies which indicate this may not be the case in Section 5, and consider the applicability of a ‘unified’ sliding relationship in Section 6.3, but focus briefly here on the differences that still remain between the regularised-Coulomb sliding relationships of Helanow and others, (2021) and Zoet and Iverson, (2020). In order to match the relationships of Zoet and Iverson, (2020) and Helanow and others, (2021) an unrealistically low friction angle is required in the equation of Zoet and Iverson, (2020) (Fig. A1), indicating that the set-up of Zoet and Iverson, (2020) produces a stronger bed than Helanow and others, (2021) under equivalent effective pressure and velocities, despite the general conception that soft beds are weaker than their hard-bed equivalents (e.g. Koellner and others,, 2019; Joughin and others,, 2009). It is possible that the strength difference is explained by (i) a natural requirement for lower subglacial water pressure in hard beds (by around 4%), which is supported by available evidence (e.g. Engelhardt and others,, 1990; Doyle and others,, 2018), (ii) non-negligible ice-rock interface friction, which is not included in Helanow and others, (2021), or (iii) unaccounted for form drag (i.e. an increase in $\boldsymbol{\mathcal{D}}_{T}$ within Eq. 8 if sliding with cavitation described in Helanow and others,, 2021, is taken as a process within $\boldsymbol{\mathcal{S}}_{T}$) occurring in many hard-bed settings which results in a view of a stronger bed.
Last, Tsai and others, (2021) take an agnostic approach to the soft- and hard- bed division, focusing instead on the region of the bed where the hydrological system is expected to be active (where $\tau_{b}$ is locally set to 0) and inactive (where $\tau_{b}$ is locally determined using power-law sliding). Under this formulation Tsai and others, (2021) find a relationship similar to Budd and others, (1984) sliding (their Eq. 6) which matches well with annual variations in velocity and moulin discharge at a land-terminating sector of the Greenland Ice Sheet. This relationship may be well-applicable in large-scale models (Section 6.3).
3.7 Frozen-thawed transitions and sub-temperate sliding
In some simpler theories both sliding and slip are considered to be negligible for frozen beds (Lliboutry,, 1966; Cuffey and Paterson,, 2010, 7.1.1), but most models either simply neglect a temperature control on sliding (including all models inverting for basal traction with present-day velocity fields in ISMIP6 experiments) or take a heuristic approach to scaling basal traction with temperature to avoid a stress singularity at the frozen thawed boundary (e.g. Fowler and Larson,, 1978; Tarasov and Peltier,, 2007; Mantelli and others,, 2019) giving
$\tau_{b}=f(u_{b},N,T_{h})$ (16)
as an adaptation of Eq. 1 where $T_{h}=T_{b}-T_{pmp}$ is the homologous temperature calculated from the pressure melting point, $T_{pmp}$, and ice temperature at the base, $T_{b}$. This corroborates with field studies finding slow (<0.2 m a-1) slip rates even with temperatures some degrees below freezing for both soft- and hard-bed settings (Hallet and others,, 1986; Echelmeyer and Zhongxiang,, 1987; Cuffey and others,, 1999; Fitzsimons and others,, 1999). The mechanical process or processes responsible for sub-temperate sliding in these instances remains uncertain (Cuffey and Paterson,, 2010, 7.2.8) but may be related to a very thin (nm) premelt layer at the ice bed (Shreve,, 1984), the inclusion of solid friction (Fowler,, 1986; Section 3.2), or sediment deformation despite freezing conditions Waller, (2001). Separately, both form drag and geometric resistance will still be expected under freezing bed conditions, though at a slower rate than under equivalent stress conditions in colder ice. This processes is framed by Weertman, (1967) as the continuation of enhanced creep despite the cessation of regelation, and matches Eq. 8 with an absent or much reduced $\boldsymbol{\mathcal{S}}_{T}$ component. Expanding sub-temperate sliding out to a grid cell under consideration also allows for the situation of both freezing and thawed areas within the same grid cell (Fig. 8), which may be a common situation in frozen-thawed transition zones (e.g. Oswald and Gogineni,, 2012; Dawson and others,, 2024). Generally, a pattern of greater basal traction in interior ice-sheet regions where a frozen bed is considered more likely is obtained in ice-sheet wide inversions (Larour and others,, 2012; Morlighem and others,, 2013).
Figure 8: Subtemperate sliding. Triangle represents area $R$ under consideration. Solid black line represents frozen-thawed transition at the bed and cooler colours represent decreasing temperature of frozen portion of bed. Sub-temperate sliding is not sufficiently well-studied to provide a definite spatial range, but speculatively this may range from 100 m - 10 km. Adapted from Wilch and Hughes, (2000).
4 Subglacial Hydrology
Subglacial hydrology and basal sliding are intimately linked via close feedbacks in many glacial settings, with water pressure influencing sliding velocity via the effective pressure $N$ and sliding velocity in turn influencing $N$ as changes in cavitation affect hydraulic conductivity and drainage system connectivity (e.g. Schoof,, 2010). There are too many facets of this topic to cover comprehensively here (reviews include: Irvine-Fynn and others,, 2011; Ashmore and Bingham,, 2014; Chu,, 2014; Siegert and others,, 2016, Nienow and others,, 2017; Davison and others,, 2019), so we focus briefly on the importance of a spatio-temporally varying subglacial hydrology system, without considering the mechanisms themselves in detail.
In most numerical ice sheet and glacier models, subglacial hydrology is either absent or is parametrised in a simple fashion through the effective pressure $N$ term in the sliding relationship (Table. 1, Fig. 9) – with at least three materially different ways of including $N$ in Eq. 1 (Table 1). While in the physical sense the meaning of $N$ as subglacial water pressure subtracted from the overburden pressure at a point in time and space is very clear, its meaning, and ways of quantifying a representative $N$ at a given scale are not straight forward in a modelling sense. Subglacial hydrological environments are very spatially diverse even at scales of tens of metres, with various terms (e.g. ‘channelised; distributed; partly-connected’) referring to various end members in a continuum of characteristics (e.g. Hubbard and Nienow,, 1997; Hoffman and others,, 2016; Rada and Schoof,, 2018). Borehole observations and investigations of deglaciated terrains (Fig. 9) evidence this complexity, with widely varying $p_{w}$ time series observed in adjacent boreholes (Doyle and others,, 2018; Rada and Schoof,, 2018; Doyle and others,, 2021) capturing distinctly separate components of subglacial hydrology, consistent with theoretical studies finding low-pressure conduits adjacent to higher-pressure regions with reduced hydrological connectivity (Engelhardt and Kamb,, 1997; Hewitt and others,, 2018). Therefore, while the influence of $N$ is fairly well understood for specific sliding sub-processes in isolation – for example, cavitation over a relatively flat bedrock area (Helanow and others,, 2021), and sediment shearing in a laboratory setting (e.g. Iverson and others,, 1998) or through a numerical continuum model (e.g. Damsgaard and others,, 2020) – it remains to fully quantify the influence of a spatially variable $N$ within a given model grid cell. We are so far only aware of the study of Tsai and others, (2021) which locally sets $\tau_{b}$ to zero over an area considered to be comprised of cavities, or elsewise calculates it following a power-law relationship, resulting in a relationship that overall shares similarities with the Budd and Jacka, (1989) sliding relationship (Table 1).
Beyond this, it is challenging to know exactly how to relate $N$ to other parameters and variables within a standard sliding relationship. In our framework (Section 2), varying $N$ could be incorporated within the parameter $P_{R_{i}}$ in Eq. 9 or power-law equivalent, though calculating its actual spatial distribution would still remain a major challenge. Most extant cases implement a simple linear multiplication (e.g. Choi and others,, 2022), but what, for example, if the traction within a region with low $N$ for a large proportion of its area was actually controlled by a small prominent pinning-point with high $N$?
Moreover, subglacial drainage systems are also observed to evolve diurnally and seasonally in response to supraglacial melt water inputs (Hubbard and Nienow,, 1997), even beneath ice ca. 1 km thick (Chandler and others,, 2013). This evolution in turn leads to diurnal and seasonal variability in sliding (Willis,, 1995; Zwally and others,, 2002; Bartholomew and others,, 2011; Nienow and others,, 2017). Such temporal variability in $N$ is by design not captured by ice sheet or glacier models employing temporally constant parameters for the basal traction relationship. However, despite the clear importance of subglacial hydrology as a control on sliding at sub-annual scales however, and with the exception of surging glaciers (Sevestre and Benn,, 2015), its overall influence on long-term glacier and ice sheet evolution remains debated, particularly with respect to future changes of the Greenland Ice Sheet (e.g. Tedstone and others,, 2015; Doyle and others,, 2015; Beckmann and Winkelmann,, 2023), with major and sustained changes in glacier and ice-sheet velocity fields typically linked to geometry changes rather than intrinsic changes in the hydrological system (e.g. Vincent and others,, 2012; Habermann and others,, 2013; Moon and others,, 2014; Cook and others,, 2016; Seroussi and others,, 2017; Catania and others,, 2018; Dehecq and others,, 2018; Gimbert and others,, 2021; Frank and others,, 2022). However, large-scale observations do indicate that major changes in the hydrological system – such as a switch from marine-terminating to land-terminating outlet glaciers, or an advancement of the ablation zone – may influence large-scale geometry and velocity patterns over a multi-annual timescale (Maier and others,, 2022). Subglacial water pressure also plays a crucial control at marine-terminating outlet glaciers grounded below sea level, which we do not cover in detail here, where very low inferred basal traction values and subglacial water pressure very close to overburden make systems particularly susceptible to changes in $N$ (e.g. Katz and Worster,, 2010; Stearns and Van Der Veen,, 2018; Minchew and others,, 2019; Bradley and Hewitt,, 2024).
Figure 9: Subglacial hydrology as observed in the field or represented in an ice flow model element. a Cavities (pale blue), channels (dark blue lines) and striation flow markers from fieldwork at Blackfoot Glacier (Walder and Hallet,, 1979). b Simplification to a single mesh element.
5 Heuristic studies
A separate approach to investigating glacier sliding is to use large-scale remotely sensed or field data to heuristically (i.e. through process-agnostic iterative experimentation and adjustment) determine the form of the sliding relationship, which may then be compared to theoretical expectations (Section 3, Table 1). A time series with significant variation in surface geometry is best suited to the decadal time scales over which mass loss is most often considered (Gillet-Chaulet and others,, 2016; Gilbert and others,, 2023), but several other variations have been exploited (Table 2). Considered in their entirety, these studies indicate that variation in the form of the sliding relationship across variable glacier settings is not readily captured by a single tunable coefficient, $C$, obtained through a standard inversion procedure. For example, Gimbert and others, (2021) and Gilbert and others, (2023) find $m=3.1\pm 0.3$ in power-law sliding for Argentière Glacier in the French Alps using a combination of surface velocity changes, length changes, and direct sliding measurements while Gillet-Chaulet and others, (2016) find $m\leq 5$ and up to 50 well reproduces surface velocities at Pine Island Glacier, Antarctica. Maier and others, (2021) meanwhile indicate that a single $m$ value, also for power-law sliding, cannot be applied to the entire Greenland Ice Sheet with, for example, an $m$ of 8 in northeast Greenland contrasting an $m$ of 4 in central southwest Greenland. Note that in standard applications using power-law sliding, $C$ is the subject of the inversion and varies in space, while $m$ is kept constant.
Setting Variation Study $m$
Argentière, French Alps – valley glacier Length (velocity) over 100 (15) years to $\sim$ 2020 Gilbert and others, (2023) 3
Hofsjökull, Iceland – ice cap§ Summer-winter velocity comparison Minchew and others, (2016) $\infty$/plastic
Greenland Ice Sheet – ice sheet Single velocity field averaged 2005-2015 Maier and others, (2021) 3-10
Northwest Greenland – outlet glaciers Forward modelling of mass loss 2007-2018 Choi and others, (2022) 6 in modified Budd
Northeast Greenland Ice Stream – ice stream Forward modelling of mass change 2006-2021 Khan and others, (2022) Regularised-Coulomb
Western Greenland – land-terminating outlet glacier Annual velocity and hydrology variations 2006-2013 Tsai and others, (2021) 4 in modified Budd
Pine Island Glacier, Antarctica – outlet glacier Geometry 2003-2008 Joughin and others, (2010) mix of 3 and $\infty$/plastic
Pine Island Glacier, Antarctica – outlet glacier Velocity in 1996 and 2007-2010 Gillet-Chaulet and others, (2016) $\geq$5
Pine Island Glacier, Antarctica – outlet glacier Forward modelling of velocity field and geometry 2002-2017 Jouvet and Huss, (2019) Regularised-Coulomb
Rutford Ice Stream, Antarctica – ice stream Displacement over fortnightly tidal cycles Gudmundsson, (2011) 3
High Mountain Asia – valley glaciers Velocity fields 2000-2017 Dehecq and others, (2018) 4
West Greenland – outlet glacier Front position from 1985-2019 Jager and others, (2024) Regularised-Coulomb
Table 2: Heuristic studies on sliding parameters. indicates largely hard bed, §
indicates largely soft bed. Absence of a superscript indicates ambiguous or mixed conditions. Notes: Gudmundsson, (2011) did not test $m$ values above 3. Dehecq and others, (2018) combine ice deformation and sliding in a 1D model. Choi and others, (2022) also include the influence of $N$, defined as ice pressure above hydrostatic equilibrium with a sheet perfectly connected to the ocean, finding best fitting with $N$ included.
Such studies offer a testing ground for proving the efficacy of a given sliding relationship, but it is important to note that a match to a given theory does not on its own demonstrate that the physical processes that build it up are actually occurring in the study location. This is straightforward in the case of $m\leq 5$ from Gillet-Chaulet and others, (2016) at Pine Island Glacier. No existing theory postulates this relationship (Section 3) so the utility of $m\leq 5$ in power-law sliding comes either from the fact that it emulates regularised-Coulomb sliding that does have a large body of theory behind it, or it represents a hitherto unappreciated set of processes not presently incorporated into a process-based sliding relationship. Meanwhile, it is difficult to interpret the findings of Minchew and others, (2016) from a setting known to have significant till coverage (Björnsson and others,, 2003) as anything other than plastic deformation of till. In contrast, Gimbert and others, (2021) and Gilbert and others, (2023) suggest that their findings of $m\sim 3$ reflects only the ice deformation component of the original Weertman, (1957) paper without the regelation component where $m=1$, neglecting the fact that the original Weertman theory is somewhat dependent upon a combination of both regelation and ice deformation, and not making an allowance for the slip component actively observed beneath Argentière (Vincent and Moreau,, 2016; Gimbert and others,, 2021). Maier and others, (2021) also invoke Weertman-type hard-bed physics as justification for lower values of $m$, despite the fundamental issues underlying the theory (Section 3.2) and the fact that no theory building upon Weertman, (1957), or the original study, presently provide a physical justification for $n=4$ or above.
Taken collectively, heuristic studies emphasise a heterogeneity between settings that is not easily reconciled with a single sliding relationship controlled by a single tunable parameter. However, they also present perhaps the most definitive method of accurately determining the optimum sliding relationship for a given modelling setting. A standardised methodology that can be efficiently applied to varying settings utilizing increasingly long datasets could therefore be of use in transparently tailoring sliding relationships for given modelling applications.
6 Discussion
6.1 Applying an inner-outer flow framework to multiple sub-processes: bridging hard and soft beds and highlighting the importance of topographic roughness
We can separate the sub-processes covered in Section 3 that deal with continuous, rather than stick-slip, motion into the three categories defined for Eq. 8. All soft-bed processes and hard-bed processes over macroscopically flat surfaces (Sections 3.1, 3.2) can be defined as slip processes (within $\boldsymbol{\mathcal{S}}_{T}$). All processes concerning viscous deformation of ice (be it cold, the basal ice layer, or temperate, Sections 3.3, 3.4) can be defined as viscous deformation processes (within $\boldsymbol{\mathcal{D}}_{T}$). Hard-bed sliding processes involving ice deformation straddle slip, deformation, and geometric resistance. The most precise way to include these would be to deconstruct their components (i.e. free slip at the ice-bed interface and over cavities, geometric resistance from obstacles, enhanced deformation) and reimpose them as segregated sub-processes, but this would then place the onus of calculating their evolution with increasing $u_{b}$ on each usage rather than allowing for it to be more simply sub-parametrised. This creates particular complications for cavitation, which is very computationally expensive to simulate (Helanow and others,, 2021). For the purposes of this discussion we therefore treat ostensibly ‘hard-bedded’ power-law and regularised-Coulomb theories (Section 3.2) as slip processes within $\boldsymbol{\mathcal{S}}_{T}$. This involves the explicit assumption that the region $R$ is then considered as sufficiently large so as to accommodate topographic features with wavelengths >25 m, if present.
The relative balance of $\boldsymbol{\mathcal{S}}_{T}$, $\boldsymbol{\mathcal{D}}_{T}$, and $\boldsymbol{\mathcal{G}}$ within Eq. 8 then allows us to make assessments of their influence on the overall traction at the inner-outer boundary for a given setting at the scale of ice-sheet or glacier model discretisation. If, as appears reasonable for relatively smooth isolated bed sub-sections (Section 3.6), regularised-Coulomb can represent slip over both soft and hard beds in the presence of abundant meltwater, the term $\boldsymbol{\mathcal{S}}_{T}$ becomes agnostic over whether the bed is soft or hard. Then, if the region is relatively planar and $b_{io}-b_{b}$ is relatively small, $\boldsymbol{\mathcal{S}}_{T}$ dominates and it is most likely that a regularised-Coulomb relationship emerges. However, if $\boldsymbol{\mathcal{D}}_{T}$ (determined as a power-law, Eq. 10) and $\boldsymbol{\mathcal{S}}_{T}$ (related to increasing velocity, obstacle prominence, and obstacle frequency, Eq. 13) dominate, then the traction for the given region will be better represented through a power-law. This implies that bed roughness, not bed type is the crucial control on the form of the sliding relationship.
The case of frozen-bed sliding would furthermore shut off, or greatly reduce the importance of $\boldsymbol{\mathcal{S}}_{T}$, resulting in power-law sliding regardless of the roughness characteristics of the region in consideration.
6.2 Maximum cavitation sizes, and revisiting Iken’s bound
[…]
If large topographic obstacles are present, such as those in Fig. 10, but large-scale cavities are not concomitant (Fig. A3), then the given grid cell area may come to be dominated by topographic roughness rather than having cavitation essentially smooth over the ice-bed interface (Fig. A3). As discussed in Section 6.1, the stage is then set for the traction defined at the inner-outer boundary to be dominated by resistance related to viscous deformation ($\boldsymbol{\mathcal{D}}$) and bed geometry ($\boldsymbol{\mathcal{G}}$), rather than a bounded basal traction relationship overall. Nonetheless, while we can place common-sense limits on cavity size in deglaciated areas, a full understanding of the scale distribution of cavities, and its bearing upon Iken’s bound, is an important future research goal.
In addition, the term $\mathrm{tan}(\beta)$ in Iken’s bound (Eq. 14) will tend towards infinity as slopes opposing the prevailing approach perpendicular , and it is therefore no longer a bound in the situation of steeply oriented cliff faces, as seen in Fig. 10 and in many rough deglaciated landscapes. We suggest that the unsuitability of Iken’s bound in regions featuring steep slopes, coupled with the importance of large topographic obstacles in increasing the $\boldsymbol{\mathcal{D}}_{T}$ and $\boldsymbol{\mathcal{G}}$ terms in Eq. (8) at larger scales may explain the continued efficacy of a power-law sliding relationships in ice-sheet and glacier settings (Table 2).
Figure 10: Rough topography near Hardangerfjord, western Norway. Hut location is 60.4526 N, 6.1406 E, inset shows ArcticDEM topography for the valley between the hut and the small glacier in the distance with no vertical exaggeration (Porter and others,, 2018). Palaeo flow direction at around 20 kyr ago available from e.g. Mangerud and others, (2019); Jungdal-Olesen and others, (2024). Numerous very steep surfaces opposing the direction of palaeo ice sheet motion can be seen at both a micro and macro scale. Photo taken by RL in late September 2024.
6.3 Does a unified basal sliding relationship exist?
A defining objective of much research into sliding is to uncover a ‘universal’ sliding relationship which can provide an accurate description of sliding processes across all settings, scales, and time periods. Generally, the implicit assumption is that this will be controlled by a single tunable coefficient (henceforth single coefficient) and/or calculated subglacial water pressure (e.g. Minchew and Joughin,, 2020b). Unfortunately, while it may be possible to describe glacier sliding by incorporating a multitude of stress contributions (Eq. 8) the complex nature of myriad sliding sub-processes and their scale and setting dependence (Section 3) likely precludes a single-coefficient relationship applicable to all settings – an outlook supported by the diversity of optimum single-coefficient relationships in Table 2. Instead, the studies reviewed here and Eq. 8 point towards a compound relationship that bears closer resemblance to descriptions of ice strain comprising several deformation mechanisms (e.g. Eq. 3 of Goldsby and Kohlstedt,, 2001).
Nonetheless, we may be able to talk about the general nature of such a compound relationship if, as in Section 6.1, we take regularised-Coulomb (Eq. 9) as the only constitutive relationship for $\boldsymbol{\mathcal{S}}_{T}$, and then take Eq. 10 as the only constitutive relationship for $\boldsymbol{\mathcal{D}}_{T}$. This allows us to take $\boldsymbol{\mathcal{S}}_{T}$ and $\boldsymbol{\mathcal{D}}_{T}$ as single terms, rather than sums as
[…]
and
[…]
where the subscript $\,{}_{Ta}$ refers to total aggregate $P_{R_{r}}$, and ${u_{t}}_{R_{r}}$, ${m_{R_{r}}}$, $n_{R_{r}}$, and $K_{R_{r}}$ are the representative parameters for the given region. $P_{R_{r}}$ and $K_{R_{r}}$ can be defined as e.g.
$P_{R_{r}}(x,y)=\frac{1}{A_{R}}\sum_{i\in\{\alpha,\beta,\gamma,...\}}I_{R_{i}}(x,y)P_{R_{i}}A_{R_{i}}\,$ (19)
for $P_{R_{r}}$ where $A_{R_{i}}$ is the area of each sub region. $m_{R_{i}}$ and $n_{R_{i}}$ cannot be calculated as a weighted average, reducing the set to a single power is a reasonable approximation in keeping with the use of $n=3$ over all glacier and ice sheet settings. Nonetheless, careful consideration should be given to $n_{R_{i}}$ in particular, given its potential dependence on whether ice in the inner layer is temperate, or has a high sediment content (Section 3.3, 3.4). Keeping the definition for $\boldsymbol{\mathcal{G}}$, but integrating over $R$ gives
[…]
where $\boldsymbol{\mathcal{G}}_{a}$ refers to the aggregate geometric resistance.
If we assume that each $\boldsymbol{\mathcal{G}}_{a}$, $\boldsymbol{\mathcal{S}}_{T_{a}}$, $\boldsymbol{\mathcal{D}}_{T_{a}}$ then defines a representative traction component at each coordinate following a moving window average, rather than for within a specific region, and with an $R$ size that at least approaches a representative surface element (Section 1), we can set
$\boldsymbol{\tau}_{b}=\boldsymbol{\mathcal{S}}_{Ta}+\boldsymbol{\mathcal{D}}_{Ta}+\boldsymbol{\mathcal{G}}_{a}$ (21)
which can be controlled by a total of 7 parameters (though dependent on the exact nature of the function used to describe $\boldsymbol{\mathcal{G}}_{a}$). Further, in the condition that $\boldsymbol{\mathcal{G}}_{a}$ does not dominate $\boldsymbol{\tau}_{b}$ in a complicated manner (or in fact follows a power-law relationship similar to $\boldsymbol{\mathcal{D}}_{T_{a}}$), and following the other assumptions in this section, we can say that $\boldsymbol{\tau}_{b}$ should always be defined as no more than the sum of a regularised-Coulomb and power-law relationship. For the one-dimensional case aligning with Table 1 that is
$\tau_{b}=f_{rC}(u_{b})+f_{pl}(u_{b})$ (22)
where $f_{rC}(u_{b})$ represents regularised-Coulomb sliding of the form Eq. 17 and $f_{pl}(u_{b})$ is a power-law of the form Eq. 18. Effective pressure may be implicitly included in Eqs. 21, 22 through parameters in $\boldsymbol{\mathcal{S}}_{Ta}$ and $f_{rC}(u_{b})$ but is presently missing as an explicit parameter here. From the perspective of the inner-outer flow boundary, the relative contribution of $\boldsymbol{u}_{b}$ to $\boldsymbol{u}_{s}$ and $\dot{\boldsymbol{\epsilon}}$ respectively will also be a function of landscape, ice rheology, and parameters for slip at the ice-bed interface. We note $f_{pl}(u_{b})$ will not be directly related to $N$ and should therefore remain important in areas with high $\dot{\boldsymbol{\epsilon}}$ within the inner flow at both low and high water pressures. Improved understanding of $\boldsymbol{\mathcal{G}}_{a}$ would allow improved confidence regarding the applicability of Eq. 22.
However, as standard inversion procedures are not well-suited to determining multiple parameters simultaneously (Morlighem and others,, 2010) and as a compound sliding relationship would complicate implementation, the choice between a compound friction law (e.g. Eq. 21) and single-term equation from Table 1 becomes a compromise between a more realistic representation of friction and a less practical numerical implementation. It may therefore be more appropriate to select a flexible yet simple equation as a sliding relationship which is capable of representing a wide range of behaviours.
6.4 Threshold velocities for regularised-Coulomb sliding
In power-law sliding, two parameters ($C$ and $m$) control the relationship, but even in the simplest regularised-Coulomb relationship at least one further parameter is required ($u_{t}$ in our usage in Table 1) to set the transition point to bounded traction. Laboratory experiments over sediment indicate a transition to bounded traction occurs at around 50 m a-1 (Zoet and Iverson,, 2020) in a controlled setting, while the transition for cavitation over hard beds at small scales (<25 m) is more variable, and possibly as low as 5 m a-1 (Helanow and others,, 2021). Given that uniformly smooth soft beds of consistent material properties, or hard beds of consistent geometry, are somewhat unlikely, it becomes complicated to suggest a reasonable physically based $u_{t}$ value that can be universally applied. For example, Joughin and others, (2019) set a threshold velocity of 300 m a-1 for Pine Island Glacier, Antarctica which sets a clear spatial boundary for the region of bounded traction, but without process-based reasoning for the threshold velocity value. This may be appropriate for the Pine Island setting, but it’s not clear if a 300 m a-1 limit would be equally effective in, for example, rough regions of the Greenland Ice Sheet. As bounded traction significantly increases the sensitivity of outlet glaciers to climate forcing (Ritz and others,, 2015; Tsai and others,, 2015; Brondex and others,, 2019), further work determining the factors controlling the threshold velocity (Trevers and others,, 2024) is necessary. We further note that in marine-terminating settings it may be difficult to distinguish whether the $u_{t}$ value used results in a good match with observations because of an intrinsic switch in processes beyond the threshold velocity, or as a result of the increasing influence of the ocean on subglacial hydrology, with ice velocity acting as a proxy for proximity to the coast.
6.5 Which basal sliding relationship to use?
No single extant sliding relationship accounts for the full breadth of processes or observations that comprise glacier and ice-sheet sliding (Sections 3, 5), making it challenging to recommend a single-tunable-coefficient sliding relationship. Pending future investigations, we make the following suggestions based primarily on heuristic studies (Section 5), whose empirically-validated nature makes them more suitable for application to glacier and ice-sheet models.
We suggest that power-law sliding (or equivalently, pseudo-plastic, Weertman-type sliding), with a setting-dependent $m$, presents a reasonable first choice. Given information from the limited number of studies available (Table 2), and in broad agreement with our discussion (Sections 6.1-6.3), lower $m$ values ($\leq 4$) are then better suited for rougher settings while higher $m$ values ($\geq 5$) are better suited for regions where topography is expected or known to be extensively smooth. Therefore, while $C$ is the main object of the inversion procedure, $m$ may be assigned based on additional constraints, such as topographic roughness. While we do not doubt the efficacy of regularised-Coulomb in some settings (Table 2), it is simpler to manipulate a power-law relationship to approximate regularised-Coulomb using a high value of $m$ than vice versa. As models for glaciers or ice sheets frequently invert for traction coefficients over regions of inferred frozen beds (Goelzer and others,, 2020; Seroussi and others,, 2020), it is further likely that a regularised-Coulomb relationship would misrepresent these areas.
If hydrology is incorporated, the simplified implementation of Budd sliding as used in Choi and others, (2022) and set as the default for ISSM (Larour and others,, 2012), which is functionally the same as power-law sliding but with $N$ included as a multiplication factor is a good starting point. A simplified hydrological model as used in Choi and others, (2022) based on elevation above sea level may then be reasonable over annual timescales (Section 4), though we stress that this is an extremely pressing area for future research – as is the exact way in which $N$ should be incorporated.
6.5.1 A note on terminology
We hope that this paper provides a clear description of the inner flow sub-processes that a sliding relationship will account for under ideal circumstances, and may also ease some of the angst experienced by modelling practitioners (ourselves included) in labelling bed properties. When we invert for basal properties we are not solely inverting for ‘roughness’, ‘till friction angle’, or ‘friction’ at the ice-bed interface itself – we are likely inverting for all of these processes, and several more, simultaneously. We use ‘ideal circumstances’ here because the traction field is often one of the few tunable parameters in an ice-sheet model and so will in practice also account for so-called ‘compensating errors’ arising from incorrectly prescribed englacial temperature, ice viscosity, subglacial topography, and surface and basal mass balance (Berends and others,, 2023). Last, while we use ‘glacier sliding’ throughout this paper to remain in keeping with its common usage, but make a distinction with ‘slip’ to refer strictly to ice-bed displacement; ‘glacier basal motion’ (e.g. Waller,, 2001; Hubbard,, 2002; Law and others,, 2023) is probably a more precise term.
7 Conclusions
Glacier sliding is not a straightforward topic. Here, we have done our best to integrate existing understanding into a relatively straightforward framework. This is centred upon an inner flow (Fig. 5) comprised of potentially numerous sliding sub-processes (Fig. 4), which can be represented over a given region through Eq. 8. While it may be possible to reduce this complexity into a relatively simple aggregate equation (such as Eq. 21), we would advocate for the use of a power-law relationship, with careful selection of the $m$ parameter based on the setting, due to its flexibility in representing both regularised-Coulomb and power-law end members.
Our coverage also highlights some clear priorities for future work. Principally, we still do not have a clear methodology for linking stick-slip sliding with continuous sliding relationships, and major questions remain regarding how best to incorporate $N$ within sliding relationships, and how it may be expected to change at the decadal to centennial timescales relevant to projections of ice-sheet and glacier mass loss. The landscape factors influencing the balance of slip, deformation, and geometric resistance also remain unexplored in a quantitative manner, as does the importance and formation, or lack thereof, of subglacial cavitation at large scales. Overall, we hope that the framework presented here allows for a transparent approach to understanding glacier sliding, applying glacier sliding relationships to numerical models, and for directing future research.
Appendix A1 Reasoning behind line placements in Fig. 2.
Form drag, sediment from Zoet and Iverson, (2020) and Minchew and Joughin, (2020b) based on the smallest clast sizes used in Zoet and Iverson, (2020) but assumed to continue to sub-milimetre scales (the ice crystal size used is milimetre-scale) and extended to a reasonable but conservative upper limit for subglacial debris size. Skin friction, sediment from Zoet and Iverson, (2020) and Minchew and Joughin, (2020b) based on the size of the shear-ring apparatus used in Zoet and Iverson, (2020) and extending to sub milimetre-scale of subglacial till sample used. Assumed to continue to sub-milimetre scales and to be valid at higher spatial scales given low topograhic variation. Weertman sliding from Weertman, (1957) with 1 cm to 10 m scale clearly stated in the text. Budd smaller-scale segment from laboratory tests in Budd and others, (1979) and larger-scale segment from application to west Antarctica in Budd and others, (1984). Nye-Kamb sliding Based on landscape used in Fig. 2 of Nye, (1970). Rate-and-state, Argentière using assumed seismic rupture length in Helmstetter and others, (2015). Cavitation theory from Lliboutry, (1968), Gagliardini and others, (2007), and Helanow and others, (2021) amongst many others. Cavitation theory is often non-dimensionalised but has not exceeded the 25 m scale used in Helanow and others, (2021). The dashed arrow is extended to 100 m to denote a plausible but untested upper limit to cavitation. Intermediate scale processes taken from the lower resolution and upper domain size used in Law and others, (2023) and extended upwards with a dashed arrow to indicate an untested upper limit for the processes described. Argentière Wheel based on the Argentière wheel experiments of Vivian and Bocquet, (1973), Gimbert and others, (2021) and Gilbert and others, (2023) amongst others where the glacier width is $\sim$300 m and 20 m is roughly double the cavity length. Skin drag from Kyrke-Smith and others, (2018) based on the lowermost resolution of radar data (40 m) to the 5 km upper limit of Bedmap2 (Fretwell and others,, 2013). Form drag, topographic from Bingham and others, (2017) and Kyrke-Smith and others, (2018) using a lower radar data resolution of 40-100 m and an upper domain area of 20 km. ’Unified’, from 1 km lower limit suggested in Tsai and others, (2021) to a plausible upper limit of 10 km. Rate-and-state, Siple Coast Taken as broad area over which seismicity is recorded at Siple Coast in Fig. 4 of Podolskiy and Walter, (2016) in contrast to single events of Helmstetter and others, (2015). Form drag, resolution/scale dependent based on arguments within this paper (Section 3.3). Typical glacier and ice sheet model resolution based on standard resolutions from detailed glacier studies to coarse resolution paleo simulations. Typical basal ice crystal size from Thorsteinsson and others, (1995) and Cook and others, (2007).
Figure A2: Schematic of a set of sliding processes operating within an inner flow. Sub-processes in cubes should be recognisable from Fig. 4. Red arrows in main panel represent variable basal slip rate and blue arrows represent form drag. Grey mesh with nodes represents model grid cells at the basal boundary.
Appendix A2 Pressure drag from a spherical obstacle
Taking a power law $\tau=M\dot{\epsilon}^{p}$ where $M$ is the consistency index, with an effective viscosity $\mu_{eff}=M\dot{\epsilon}^{p-1}$ and using $\dot{\epsilon}\sim V/L$ where $V$ is the far-field velocity and $L$ is the obstacle size gives
$\mu_{eff}\sim M\frac{V}{L}^{p-1}.$ (23)
Separately, taking the resistance, $F_{D}$, from the normal forces acting on the sphere as
$F_{D}\propto\mu_{eff}VL$ (24)
from Stoke’s law (e.g. Happel and Brenner,, 1983, 4-17.28) and substituting Eq. 23 gives $F_{D}\sim MV^{p}L^{2-p}$ which then gives Eq. 12 if we set $p=1/n$ and $M=K_{R_{i}}$.
Appendix A3 Example expansion of $\boldsymbol{\mathcal{S}_{T}}$
$\boldsymbol{\mathcal{S}}_{T}(x,y)=\begin{cases}\boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\beta}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\gamma}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\cap R_{\beta}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\cap R_{\gamma}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\beta}\cap R_{\gamma}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\cap R_{\beta}\cap R_{\gamma}.\end{cases}$ (25)
Appendix A4 Sliding relationships in geodynamics
In geodynamics problems where rate-and-state friction theory is the more common way of viewing slip, models tend to use either a free-slip between mechanical layers (e.g. Sizova and others,, 2010; Nakakuki and Mura,, 2013), a pseudo-plastic yielding based on the coefficient of frictional sliding where boundaries between layers are not made explicit but a slip horizon is created within a continuum (e.g. Tackley,, 2000; Schmalholz and Duretz,, 2015), or focus on shear-localization through a non-Newtonian rheology with grain-size evolution in an initially rheologically homogeneous media (Bercovici,, 2003 and references therein). These settings tend to have much more planar interfaces and do not feature the topographic variability that characterises some glacier beds, but provide interesting comparisons for glacier sliding theory.
Figure A3: A spatial bound to Iken’s bound? a Cavitation (blue) occurring where topography is sufficiently small adapted from Schoof, (2005). b Cavitation can occur at a sufficiently small scale (yellow highlight) but will not occur in larger wavelength depressions. z is height from baseline and x is distance along flow. The thick black line is the glacier bed. $R$ represents a possible region in consideration for the sliding relationship. This figure is drawn in two dimensions for comparison with previous studies, but the same principles apply in three dimensions with realistic topography.